OpenStudy (anonymous):

When I do the work, I get 4y = 3 +40!! What am I doing wrong? Given the equation y − 4 = 3 over 4(x + 8) in point-slope form, identify the equation of the same line in slope-intercept form. −3 over 4x + y = 10 3x – 4y = −40 y = 3 over 4x + 12 y = 3 over 4x + 10

OpenStudy (anonymous):

y=mx+c is standard slope intercept form arrange equa to compare with this

OpenStudy (anonymous):

I know that y = mx +c but my answer doesn't match any of the multiple choice questions... !!! I don't know what I'm doing wrong..

OpenStudy (anonymous):

is it y-4=(3/4)*(x+8)

OpenStudy (anonymous):

The question is y − 4 = 3/4(x + 8) put into slope intercept form.

OpenStudy (anonymous):

4y-16=(3x/4)+24 4y-16=(3x+96)/4 16y-64=3x+96

OpenStudy (anonymous):

16y=3x+160 y=(3x/16)+10 slope=3/16 yintercept=10